# Breaking the Trend: How to Avoid Cherry-Picked Signals

### v10

@@ 3 Theory from Grebenkov and Serror (2014)
Eq. 3 differs slightly when assuming that the correlations between $\xi_{i,k}$ are zero, as shown in their Eq. (43) [^12] Eq. (43) $\sqrt{\frac{Nq^{2}(1-p^{2})}{Q^{2}(1-\bm{\rho_{\epsilon}}^{2})+2Q(1-\bm{\rho_{\epsilon}})+R}}$ where $p=1-\eta$, $q=1-\lambda$, and $Q=\frac{(1-pq)\sigma^{4}}{\beta_{0}^{4}}$., but this correlation-dependent formulation remains problematic: it fails to converge as $N\to\infty$, which contradicts empirical observations, suggesting that the underlying assumption (that correlations between $\xi_{i,k}$ are zero) is unrealistic.
@@
Nevertheless, we can reasonably expect Eq. 3 to hold more precisely and perhaps even exactly—thanks to rotational invariance  [^13] We know from Grebenkov and Serror (2015) and the discussion above that Eq. 3 holds when considering coordinate transformation to the eigenvectors of the correlation matrix of $\epsilon_{i,t}$, in which the ”random” correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$ would be set to zero under the $ARP$ conditions.—under the more realistic conditions that make the ARP, and not the Markowitz portfolio, optimal.
@@
In that case where Eq. 3 holds, Eq. 2 must be adjusted to redefine $\beta_{0}$. To be consistent with the observed increase of the Sharpe ratio with the size of the universe $N$, we can, under strong approximations, expect $\beta_{0}$ to be related to $\frac{\beta}{\sqrt{\lambda\left(2-\lambda\right)}}\sqrt{\frac{N}{1+\left(N-1\right)\bm{\rho}^{2}}}$ where $\bm{\rho}^{2}$ is the average squared correlation in the universe. This expression extends Eq. 2 using Eq. (40)  [^14] Their Eq. (40) served only as an inspiration, as it was derived under the assumption of identical correlation coefficients for the two matrices (the correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$)—conditions under which the Markowitz portfolio would be optimal, but not the $ARP$. This inconsistency is offset by another one: the Eq. (40) relates only the scaling factor between the optimal Sharpe and the naïve ’1/N’ Sharpe. from Grebenkov and Serror (2015). The rescaling factor formula, $\sqrt{\frac{N}{1+(N-1)\bm{\rho}^{2}}}$, is empirically tested through the relationship between the Sharpe ratio and $N$ in Section 5.2. This rescaling factor formula also corresponds exactly to the theoretical scaling of the Sharpe ratio for a portfolio of $N$ assets with identical expected returns and volatilities, and an average pairwise correlation of $\rho^{2}$. This makes sense, as the correlation between two trend-following strategies built on underlyings correlated at $\rho$ is expected to be $\rho^{2}$. Thus, the scaling factor for ARP ultimately coincides with that obtained for the naïve ’1/N’ trend-following portfolio, which does not use the correlation matrix to determine positions.
@@
Additionally Grebenkov and Serror (2014) derived in Eq.4, the optimal smoothing parameter ($\eta_{\text{opt}}$) of the Exponential Moving Average ($\mathrm{EMA}$) applied to returns as a trend indicator for a trend-following strategy which is supposed to yield to the optimal Sharpe ratio. It is interesting to observe from Eq.4 that $\frac{1}{\eta_{\text{opt}}}$, the time scale of the signal $\mathrm{EMA}$ generating the optimal strategy is always shorter than $\frac{1}{\lambda}$, the time scale of the best predictor $\mathrm{EMA}$ (or the optimal indicator in the sub optimal optimization problem of Zakamuli and Giner (2024)) and that of the autocorrelated process of returns. Indeed, when $\beta_{0}$ increases (either due to a stronger trend or a more diversified universe), the time scale of the optimal $\mathrm{EMA}$ should be reduced. As a conclusion, theoretically the parameters of the signals should be adjusted when the universe is increased.
@@
Figure 3: Theoretical Sharpe ($\mathbf{S}$} from Eq. 3) of the trend following strategy depending on $\eta$ the parameter of the $\mathrm{EMA}$. $\beta_{0}=0.1$ and $\lambda=0.01$ for different cost of trading $\theta$, replicated from Grebenkov and Serror (2014)
@@ 4 Description of the empirical analysis
To validate Grebenkov’s theory (Eq. 3), introduced in Sec 3 and, therefore, the diffusive process in Eq. 1 which is assumed for the underlying assets, we conducted empirical backtests of the Agnostic Risk Parity (ARP) strategy introduced by Benichou at al (2017). We decided to use the $ARP$ as the projection portfolio, as it is theoretically and empirically considered to be the optimal trend-following portfolio, capable of minimizing measurement noise in the Sharpe ratio. However, we expect that the results should also hold even with lower Sharpe ratios and more noises for more conventional and naive ’1/N’ trend-following portfolios, such as those that do not manage correlations or assume an identity correlation matrix except when computing volatility used to target the size of the portfolio.
The other objective advantage of $ARP$ is that it constructs a rotationally invariant portfolio which makes the Sharpe ratio less dependent on the subjective selection of futures used to define the investment universe, as well as on discretionary choices regarding the relative number of FX, stock index, bond, and commodity futures. $ARP$ also by consequence applies trend-following to eigenvenvectors while allocating equal unconditional risk to each eigenvector of the covariance matrix. While rotational invariance is a desirable theoretical property, the equal unconditional risk allocation among eigenvectors is a more desired property for asset managers. Moreover, Valeyre (2024) demonstrated that the ARP portfolio is optimal—achieving the highest Sharpe ratio—under the assumption that the covariance of the $\xi_{i,k}$ terms in Eq. 1 follows a particular random matrix structure, which is considered realistic. Implementing the ARP portfolio is relatively straightforward and consists of normalizing the trend indicators by the inverse square root of the correlation matrix, rather than by the full inverse as in the classical Markowitz framework.
We decided to focus primarily on exponential moving averages ($\mathrm{EMA}$) as trend indicators to feed the Agnostic Risk Parity (ARP) strategy, as they are theoretically optimal according to Grebenkov and Serror (2014); Zakamuli and Giner (2024), assuming the validity of Eq. 1. Accordingly, we tested various $\mathrm{EMA}$ configurations and measured how the empirical Sharpe ratio of the resulting portfolio depends on the EMA parameter, in line with Grebenkov’s theoretical predictions. Additionally, we explored more complex trend indicators based on multiple time scales—such as the Moving Average Convergence Divergence ($\mathrm{MACD}$)—to empirically validate whether the $\mathrm{EMA}$ remains the optimal choice.
We backtested the Agnostic Risk Parity (ARP) strategy on a global universe including Futures on commodities, FX, stock indices and bonds on a period 25th of May 1990 and stops on the 7th of December 2023.
@@ 4.1 Renormalized exponential moving average indicator ($\mathrm{EMA}$)
We used only a normalized version of the $EMA$ to account for the dependence of volatility on both time and instrument. This normalization allows for meaningful comparisons of trend indicators across instruments, time periods, and even different smoothing parameters. At time $t$, the $\mathrm{EMA}$ is a vector of $N$ exponential moving averages of normalized returns, one for each of the $N$ instruments. The $EMA$ of instrument $i$ at time $t+1$ is defined by $\phi$ in Eq. 6, where $r_{i,t+1}$ denotes the return of instrument $i$ at time $t+1$. The parameter $\sigma$ represents the standard deviation of returns used to normalize them, and $\eta$ is the smoothing parameter of the $EMA$ as well as of the return normalization. $\sigma^{2}$ can be interpreted as an exponential moving average of squared returns, and the indicator $\phi$ corresponds to an exponential moving average of returns scaled by their volatility. The incrementation is on a daily basis but it could be generalized to minutes returns.
It is interesting to note that the half-life (originally introduced in nuclear physics), defined as the lookback period over which the cumulative weight of past observations reaches $50\%$, corresponds approximately to $\frac{\ln(2)}{\eta}$. That value could be compared to the half of the look-back period of the momentum indicator, which is a more standard and usual indicator in the financial industry. By contrast, the standard number of periods used in technical analysis and in financial markets for the exponential moving average, as defined in Wilder (1978); Murphy (1999), is $\frac{2}{\eta}$. This latter quantity corresponds to an average lookback period, but it is not well suited to use a weighted arithmetic average of periods. As a result, it has weaker economic interpretation and cannot be meaningfully compared with the parameters of other indicators.
@@
The indicator $\mathrm{EMA}\left(150\right)$ is defined by the $\phi_{i,t+1}\left(\eta\right)$ when $\eta=\frac{1}{150}$ is applied to every underlying $i$ and time $t$. We use $\sqrt{\eta}$ in Eq.6 so that the std of $\phi_{i,t+1}\left(\eta\right)$ is 1 theoretically if $r$ are well approximated by unautocorrelated returns so that $\phi$ is a normalized $EMA$ indicator.
@@
Inspired by $\mathrm{MACD}$, we introduced a combination of $\mathrm{EMA}$ applying the Eq.7 while determining $\omega_{1}$ so that the derivative of the sensitivity to past daily returns is at 0 at the lag 0 and replicate curves in Fig.4. The idea behind this indicator was to reduce the weight of returns at very short time scales—where mean reversion is possible—and to increase the weight of returns at longer time scales, in order to test a multi-timescale memory effect for trend-followers.
@@
The derivative at zeros yields to Eq.8.
@@ 4.3 Agnostic Risk Parity ($\mathrm{ARP}$) and naive ’1/N’ portfolio
We first estimated the correlation matrix $C$, of dimension $N\times N$, using a 750-day exponential moving average applied to weekly returns when implementing the RIE filter introduced by Bun at al (2016). The vector $\Sigma$ consists of N values representing the standard deviations estimated using a 40-day exponential moving average applied to daily returns.
Next we expressed the positions vector as a linear function of the signals ($\mathrm{EMA}$ through $\phi$ or $\mathrm{MACD}$) applying a normalisation that involves the inverse of the square root of the estimated correlation matrix and volatilities. This follows the formulation of the agnostic risk parity portfolio ($\mathrm{ARP}$}) introduced by Benichou at al (2017).
We then used the parameter $\rho=\frac{1}{20}$ for portfolio smoothing as specified in Eq.LABEL:eq:arp2 which is an very easy solution to reduce trading cost (usually, trend indicators include rules to reduce trading costs. While the $EMA$ does not explicitly aim to avoid excessive trading, smoothing the portfolio has proven to be both simple and effective. Benichou at al (2017) recommended optimization under constraints to limit trading and illiquidity and to better manage trading costs).
Finally, we applied a resizing process in Eq.13 to target a constant volatility for the final positions $\mathrm{ARP}$} which is a vector of N weights at time $t+1$. The target-volatility step is both usual and useful to get standardized and homoscedastic returns for the portfolio so that Sharpe ratio is adapted.
We applied exactly the same computation for the naive ’1/N’ approach, except that we assumed $C^{-0.5}$ in Eq. 12 to be the identity matrix. This is equivalent to managing trend-following strategies independently, underlying by underlying, while constraining the overall portfolio to maintain constant volatility. This naive approach is referred to as ‘1/N’ by Benichou at al (2017). We did not test the true Markowitz portfolio, which would simply replace $C^{-0.5}$ by $C^{-1}$, as that solution performed significantly worse. We also test the naive ’1/N’ approach derived from the binary trading rule where we replace $\phi_{t}$ by its sign.
@@ 4.4 Dataset and different simulated parameters
We tested the different indicators applying the $\mathrm{ARP}$ formula Eq.13. The different parameters are described in Tab 3 in the appendix D.
@@
(c) $\mathrm{ARP}\left(0\times 20,80,0.2\right)$
(d) $\mathrm{ARP}\left(0\times 20,80,0.4\right)$
(e) $\mathrm{ARP}\left(20,80,0.2\right)$
(f) $\mathrm{ARP}\left(20,80,0.4\right)$
Figure 4: Different sensitivities to past daily returns for indicators ($\mathrm{MACD}$, $\mathrm{EMA}$) all based on 20, 80 and 400 day time scales. we have $\mathrm{ARP}\left(0\times 20,80,0.2\right)$, $\mathrm{ARP}\left(0\times 20,80,0.4\times 400\right)$, $\mathrm{ARP}\left(20,80,0.2\times 400\right)$, $\mathrm{ARP}\left(20,80,0.4\times 400\right)$, $\mathrm{ARP}\left(80\right)$, $\mathrm{ARP}\left(400\right)$

### v11

@@ 3 Theory from Grebenkov and Serror (2014)
Eq. (3) differs slightly when assuming that the correlations between $\xi_{i,k}$ are zero, as shown in their Eq. (43) [^12] Eq. (43) $\sqrt{\frac{Nq^{2}(1-p^{2})}{Q^{2}(1-\bm{\rho_{\epsilon}}^{2})+2Q(1-\bm{\rho_{\epsilon}})+R}}$ where $p=1-\eta$, $q=1-\lambda$, and $Q=\frac{(1-pq)\sigma^{4}}{\beta_{0}^{4}}$, but this correlation-dependent formulation remains problematic: it fails to converge as $N\to\infty$, which contradicts empirical observations, suggesting that the underlying assumption (that correlations between $\xi_{i,k}$ are zero) is unrealistic.
@@
Nevertheless, we can reasonably expect Eq. (3) to hold more precisely and perhaps even exactly—thanks to rotational invariance  [^13] We know from Grebenkov and Serror (2015) and the discussion above that Eq. (3) holds when considering coordinate transformation to the eigenvectors of the correlation matrix of $\epsilon_{i,t}$, in which the ”random” correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$ would be set to zero under the $\mathrm{ARP}$ conditions.—under the more realistic conditions that make the $\mathrm{ARP}$, and not the Markowitz portfolio, optimal.
@@
In that case where Eq. (3) holds, Eq. (2) must be adjusted to redefine $\beta_{0}$. To be consistent with the observed increase of the Sharpe ratio with the size of the universe $N$, we can, under strong approximations, expect $\beta_{0}$ to be related to $\frac{\beta}{\sqrt{\lambda\left(2-\lambda\right)}}\sqrt{\frac{N}{1+\left(N-1\right)\bm{\rho}^{2}}}$ where $\bm{\rho}^{2}$ is the average squared correlation in the universe. This expression extends Eq. (2) using Eq. (40)  [^14] Their Eq. (40) served only as an inspiration, as it was derived under the assumption of identical correlation coefficients for the two matrices (the correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$)—conditions under which the Markowitz portfolio would be optimal, but not the $\mathrm{ARP}$. This inconsistency is offset by another one: the Eq. (40) relates only the scaling factor between the optimal Sharpe and the naïve ’1/N’ Sharpe. from Grebenkov and Serror (2015). The rescaling factor formula, $\sqrt{\frac{N}{1+(N-1)\bm{\rho}^{2}}}$, is empirically tested through the relationship between the Sharpe ratio and $N$ in Section 5.2. This rescaling factor formula also corresponds exactly to the theoretical scaling of the Sharpe ratio for a portfolio of $N$ assets with identical expected returns and volatilities, and an average pairwise correlation of $\bm{\rho}^{2}$. This makes sense, as the correlation between two trend-following strategies built on underlyings correlated at $\bm{\rho}$ is expected to be $\bm{\rho}^{2}$. Thus, the scaling factor for $\mathrm{ARP}$ ultimately coincides with that obtained for the naïve ’1/N’ trend-following portfolio, which does not use the correlation matrix to determine positions.
@@
Additionally Grebenkov and Serror (2014) derived in Eq. (4), the optimal smoothing parameter ($\eta_{\text{opt}}$) of the Exponential Moving Average ($\mathrm{EMA}$) applied to returns as a trend indicator for a trend-following strategy which is supposed to yield to the optimal Sharpe ratio. It is interesting to observe from Eq. (4) that $\frac{1}{\eta_{\text{opt}}}$, the time scale of the signal $\mathrm{EMA}$ generating the optimal strategy is always shorter than $\frac{1}{\lambda}$, the time scale of the best predictor $\mathrm{EMA}$ (or the optimal indicator in the sub optimal optimization problem of Zakamuli and Giner (2024)) and that of the autocorrelated process of returns. Indeed, when $\beta_{0}$ increases (either due to a stronger trend or a more diversified universe), the time scale of the optimal $\mathrm{EMA}$ should be reduced. As a conclusion, theoretically the parameters of the signals should be adjusted when the universe is increased.
@@
Figure 3: Theoretical Sharpe ($\mathbf{S}$} from Eq. (3)) of the trend-following strategy depending on $\eta$ the parameter of the $\mathrm{EMA}$. $\beta_{0}=0.1$ and $\lambda=0.01$ for different cost of trading $\theta$, replicated from Grebenkov and Serror (2014)
@@ 4 Description of the empirical analysis
To validate Grebenkov’s theory (Eq. (3)), introduced in Sec.3 and, therefore, the diffusive process in Eq. (1) which is assumed for the underlying assets, we conducted empirical backtests of the Agnostic Risk Parity ($\mathrm{ARP}$) strategy introduced by Benichou et al (2017). We decided to use the $\mathrm{ARP}$ as the projection portfolio, as it is theoretically and empirically considered to be the optimal trend-following portfolio, capable of minimizing measurement noise in the Sharpe ratio. However, we expect that the results should also hold even with lower Sharpe ratios and more noises for more conventional and naive ’1/N’ trend-following portfolios, such as those that do not manage correlations or assume an identity correlation matrix except when computing volatility used to target the size of the portfolio.
The other objective advantage of $\mathrm{ARP}$ is that it constructs a rotationally invariant portfolio which makes the Sharpe ratio less dependent on the subjective selection of futures used to define the investment universe, as well as on discretionary choices regarding the relative number of FX, stock index, bond, and commodity futures. $\mathrm{ARP}$ also by consequence applies trend-following to eigenvenvectors while allocating equal unconditional risk to each eigenvector of the covariance matrix. While rotational invariance is a desirable theoretical property, the equal unconditional risk allocation among eigenvectors is a more desired property for asset managers. Moreover, Valeyre (2024) demonstrated that the $\mathrm{ARP}$ portfolio is optimal—achieving the highest Sharpe ratio—under the assumption that the covariance of the $\xi_{i,k}$ terms in Eq. (1) follows a particular random matrix structure, which is considered realistic. Implementing the $\mathrm{ARP}$ portfolio is relatively straightforward and consists of normalizing the trend indicators by the inverse square root of the correlation matrix, rather than by the full inverse as in the classical Markowitz framework.
We decided to focus primarily on exponential moving averages ($\mathrm{EMA}$) as trend indicators to feed the Agnostic Risk Parity ($\mathrm{ARP}$) strategy, as they are theoretically optimal according to Grebenkov and Serror (2014); Zakamuli and Giner (2024), assuming the validity of Eq. (1). Accordingly, we tested various $\mathrm{EMA}$ configurations and measured how the empirical Sharpe ratio of the resulting portfolio depends on the $\mathrm{EMA}$ parameter, in line with Grebenkov’s theoretical predictions. Additionally, we explored more complex trend indicators based on multiple time scales—such as the Moving Average Convergence Divergence ($\mathrm{MACD}$)—to empirically validate whether the $\mathrm{EMA}$ remains the optimal choice.
We backtested the Agnostic Risk Parity ($\mathrm{ARP}$) strategy on a global universe including Futures on commodities, FX, stock indices and bonds on a period 25th of May 1990 and stops on the 7th of December 2023.
@@ 4.1 Renormalized exponential moving average indicator ($\mathrm{EMA}$)
We used only a normalized version of the $\mathrm{EMA}$ to account for the dependence of volatility on both time and instrument. This normalization allows for meaningful comparisons of trend indicators across instruments, time periods, and even different smoothing parameters. At time $t$, the $\mathrm{EMA}$ is a vector of $N$ exponential moving averages of normalized returns, one for each of the $N$ instruments. The $\mathrm{EMA}$ of instrument $i$ at time $t+1$ is defined by $\phi_{i,t}$ in Eq. (6), where $r_{i,t}$ denotes the return of instrument $i$ at time $t$. The parameter $\sigma_{i,t}$ represents the standard deviation of returns used to normalize them, and $\eta$ is the smoothing parameter of the $\mathrm{EMA}$ as well as of the return normalization. $\sigma^{2}$ can be interpreted as an exponential moving average of squared returns, and the indicator $\phi$ corresponds to an exponential moving average of returns scaled by their volatility. The incrementation is on a daily basis but it could be generalized to minutes returns.
It is interesting to note that the half-life (originally introduced in nuclear physics), defined as the lookback period over which the cumulative weight of past observations reaches $50\%$, corresponds approximately to $\frac{\ln(2)}{\eta}$. That value could be compared to the half of the look-back period of the momentum indicator, which is a more standard and usual indicator in the financial industry. By contrast, the standard number of periods used in technical analysis and in financial markets for the exponential moving average, as defined in Wilder (1978); Murphy (1999), is $\frac{2}{\eta}$. This latter quantity corresponds to an average lookback period, but it is not well suited to use a weighted arithmetic average of periods, and may be a weighted harmonic average would be better here. As a result, it has weaker economic interpretation and cannot be meaningfully compared with the parameters of other indicators.
@@
The indicator $\mathrm{EMA}\left(150\right)$ is defined by the $\phi_{i,t+1}\left(\eta\right)$ when $\eta=\frac{1}{150}$ is applied to every underlying $i$ and time $t$. We use $\sqrt{\eta}$ in Eq. (6) so that the std of $\phi_{i,t+1}\left(\eta\right)$ is 1 theoretically if $r$ are well approximated by unautocorrelated returns so that $\phi$ is a normalized $\mathrm{EMA}$ indicator.
@@
Inspired by $\mathrm{MACD}$, we introduced a combination of $\mathrm{EMA}$ applying the Eq. (7) while determining $\omega_{1}$ so that the derivative of the sensitivity to past daily returns is at 0 at the lag 0 and replicate curves in Fig.4. The idea behind this indicator was to reduce the weight of returns at very short time scales—where mean reversion is possible—and to increase the weight of returns at longer time scales, in order to test a multi-timescale memory effect for trend-followers.
@@
The derivative at zeros yields to Eq. (8).
@@ 4.3 Agnostic Risk Parity ($\mathrm{ARP}$) and naive ’1/N’ portfolio
We first estimated the correlation matrix $C$, of dimension $N\times N$, using a 750-day exponential moving average applied to weekly returns when implementing the RIE filter introduced by Bun et al (2016). The vector $\Sigma$ consists of N values representing the standard deviations estimated using a 40-day exponential moving average applied to daily returns.
Next we expressed the positions vector as a linear function of the signals ($\mathrm{EMA}$ through $\phi$ or $\mathrm{MACD}$) applying a normalisation that involves the inverse of the square root of the estimated correlation matrix and volatilities. This follows the formulation of the agnostic risk parity portfolio ($\mathrm{ARP}$}) introduced by Benichou et al (2017).
We then used the parameter $\rho=\frac{1}{20}$ for portfolio smoothing as specified in Eq. (LABEL:eq:arp2) which is an very easy solution to reduce trading cost (usually, trend indicators include rules to reduce trading costs. While the $\mathrm{EMA}$ does not explicitly aim to avoid excessive trading, smoothing the portfolio has proven to be both simple and effective. Benichou et al (2017) recommended optimization under constraints to limit trading and illiquidity and to better manage trading costs).
Finally, we applied a resizing process in Eq. (13) to target a constant volatility for the final positions $\mathrm{ARP}$} which is a vector of N weights at time $t+1$. The target-volatility step is both usual and useful to get standardized and homoscedastic returns for the portfolio so that Sharpe ratio is adapted.
We applied exactly the same computation for the naive ’1/N’ approach, except that we assumed $C^{-0.5}$ in Eq. (12) to be the identity matrix. This is equivalent to managing trend-following strategies independently, underlying by underlying, while constraining the overall portfolio to maintain constant volatility. This naive approach is referred to as ‘1/N’ by Benichou et al (2017). We did not test the true Markowitz portfolio, which would simply replace $C^{-0.5}$ by $C^{-1}$, as that solution performed significantly worse. We also test the naive ’1/N’ approach derived from the binary trading rule where we replace $\phi_{t}$ by its sign.
@@ 4.4 Dataset and different simulated parameters
We tested the different indicators applying the $\mathrm{ARP}$ formula Eq. (13). The different parameters are described in Tab.3 in the appendix D.
@@
(c) $\mathrm{ARP}\left(0\times 20,80,0.2\times 400\right)$
(d) $\mathrm{ARP}\left(0\times 20,80,0.4\times 400\right)$
(e) $\mathrm{ARP}\left(20,80,0.2\times 400\right)$
(f) $\mathrm{ARP}\left(20,80,0.4\times 400\right)$
Figure 4: Different sensitivities to past daily returns for indicators ($\mathrm{MACD}$, $\mathrm{EMA}$) all based on 20, 80 and 400 day time scales. we have $\mathrm{ARP}\left(0\times 20,80,0.2\times 400\right)$, $\mathrm{ARP}\left(0\times 20,80,0.4\times 400\right)$, $\mathrm{ARP}\left(20,80,0.2\times 400\right)$, $\mathrm{ARP}\left(20,80,0.4\times 400\right)$, $\mathrm{ARP}\left(80\right)$, $\mathrm{ARP}\left(400\right)$